Italian Queue 2014 Prev

Sep 20-21

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: MODS

Summary

At the beginning of the night the sky is really cloudy, humidity 93.8%, dewpoint 9.3\xB0C, temperature 7.9\xB0C, the TO is checking the conditions

Data Summary

Project_RUN Objects Filter ExpTime Notes
Calibrations        
FusiPecci O MCGC9 dual grating 1 h COMPLETED
FusiPecci P MCGC8 dual grating 30 m COMPLETED
Standard Feige110 dual grating 5 m  

Details

02:20 execMODS modsWake.pro

02.26 execMODS sieveSnap.pro

02:55 At the moment it is not possible to open the dome, so we perform some MODS calibrations
  • Lardo_A1 execMODS mods.3.M71.cal mods1b.20140921.0002-0006
  • Lardo_A9 execMODS M15_mos.cal mods1b.20140921.0007-0011
  • Saracco_F execMODS mods.5.xllsm1.cal mods1r.20140921.0002-0013
  • Saracco_E execMODS mods.6.XLSSm2.cal mods1r.20140921.0014-0025

04:35 The TO opens the dome

04:46 Slewing to FusiPecci O acqMODS MCGC9_ls.acq mods1r.20140921.0026-0028
  • seeing 0.9''
  • modsAlign -l -B mods1r.20140921.0026-0027 the -B in modsAlign allows to change the scale in the modsAlign SAOImage window
  • during the acquisition MODS sent an error message about dewar pressure; in source-slit image (mods1r.20140921.0028) the source appears less bright probably due to some clouds passing

05:06 execMODS MCGC9_ls_04.obs mods1r.20140921.0029-0031 mods1b.20140920.0012-0014 the last images have only 500s of exposure
  • seeing 0.9''
  • clouds coming, we stop the acquisition, then resume it when they are passed
  • seeing 0.9''
  • we pause the acquisition, seeing 1.2'', then we abort it, guiding star lost
  • the last images mods1r.20140921.0031 and mods1b.20140920.0014 have only 500s of exposure
  • we decided to take two more spectra because during the exposition there were clouds and the last exposure was stopped abruptly

06:40 again FusiPecci O due to clouds we have to acquire again acqMODS MCGC9_ls.acq mods1r.20140921.0032-0034
  • We have lost the guide star so we have to acquire again
  • modsAlign -l -B mods1r.20140921.0032-0033

06:49 execMODS MCGC9_ls_04.obs mods1r.20140921.0035-0036  mods1b.20140921.0015-0017 we have taken two images each channel, the last blue image (mods1b.20140921.0017) is to be discarded
  • seeing 0.9''

07:35 slewing to FusiPecciP acqMODS MCGC8_ls.acq mods1r.20140921.0037
  • we recover the script with the offset found yesterday so not to have to align, through slit image only (mods1r.20140921.0037)
  • script modified so to acquire only two images (1800s)

07:47 execMODS MCGC8_ls_05.obs mods1r.20140921.0038-0039   mods1b.20140921.0018-0019
  • seeing 0.9''

08:22 slewing to standard Feige110 (script modified to add also red grating configuration) acqMODS feige110.acq mods1r.20140921.0040
  • we recover the script with the offset found yesterday so not to have to align, but it doesn't work, the star is not in the slit (mods1r.20140921.0040) so we have to acquire again an do the alignment
  • acqMODS feige110.acq mods1r.20140921.0041

08:31 modsAlign - r mods1r.20140921.0041

08:35 problems with the guiding star so the program automatically changed to another guiding star, causing the pointing to be offset of 20''. we have to do again acquisition.

08:37 slewing again to Feige110 acqMODS feige110.acq mods1r.20140921.0042
  • modsAlign -r mods1r.20140921.0042

08:42 execMODS Feige110.obs mods1r.20140921.0044-0046  mods1b.20140921.0020-0022
  • seeing 1''
  • we have performed only the dual grating part of the script because the TO must close the dome for the humidity

08:54 The dome is wet, the TO closes

11:07 doBias mods1r.20140921.0047-0056 mods1b.20140921.0023-0032
Topic revision: r7 - 10 Nov 2014, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback